OpenStudy (anonymous):

3/((x-8)(x-3))=1/(x-8)+8(x-3)

OpenStudy (saifoo.khan):

\[\frac{3}{(x-8)(x-3)}=\frac{1}{x-8}+\frac{8}{x-3}\]

OpenStudy (saifoo.khan):

\[\frac{3}{(x-8)(x-3)}-\frac{1}{x-8}-\frac{8}{x-3}=0\]

OpenStudy (saifoo.khan):

take the lcm = (x-8)(x-3)

OpenStudy (saifoo.khan):

And solve.! ^_^

OpenStudy (anonymous):

i cant

OpenStudy (saifoo.khan):

lol.

OpenStudy (anonymous):

help?

OpenStudy (saifoo.khan):

yes, x = 70/9